A rope is 53 m long.divide it intotwo so thatone piece is 17m shorter than the other.find the measure of each piece
Answer by Cromlix(4381)

Hi there,
One piece = 'x'
theother piece = 'x - 17'
x + (x - 17) = 53
2x - 17 = 53
2x = 53 + 17
2x = 70
x = 35
One piece = 35m long
the other piece = 18m long (x - 17)
Hope this helps:-)